A 200 glider and a 600 g glider are coupled together and move along a frictionless air track at 50 cm/s. An explosive charge on one of the gliders separates them, causing the 200 g glider to stop. What is the subsequent velocity of the 600 g glider?



1
Expert's answer
2021-03-23T08:15:53-0400

"(m_1+m_2)v=m_1\cdot 0+m_2u,"

"u=\frac{m_1+m_2}{m_2}v,"

"u=\frac{200+600}{600}\cdot 50\approx67~\frac{cm}{s}."